OpenStudy (anonymous):

I WILL GIVE MEDAL FOR ANSWER. What is the exact value of the expression square root of 18 plus square root of 48 minus square root of 50? Simplify completely.

OpenStudy (anonymous):

\[\eqalign{ \sqrt{18}+\sqrt{48}-\sqrt{50}&=\sqrt{9\times2}-\sqrt{25\times2}+\sqrt{48} \\ &=3\sqrt{2}-5\sqrt{2}+\sqrt{48} \\ &=\sqrt{48}-2\sqrt{2} \\ &=2\sqrt{12}-2\sqrt{2} \\ &=2(\sqrt{12}-\sqrt{2}) \\ &=2(\sqrt{6}\sqrt{2}-\sqrt{2}) \\ &=2[\sqrt{2}(\sqrt{6}-1)] \\ &=(2)^{1.5}(\sqrt{6}-1) \\ &=(2)^{1.5}(\sqrt{2}\sqrt{3}-1) \\ &=2^2\sqrt{3}-1 \\ &=4\sqrt{3}-1 }\] ...That's as far as I got haha

OpenStudy (anonymous):

wait i have multiple choice answers

OpenStudy (anonymous):

...That might help haha

OpenStudy (anonymous):

A.sqrt(16) B.-14sqrt(2)+16sqrt(3) C.-2sqrt(2)+4sqrt(3) D.-sqrt(8)+sqrt(48)

OpenStudy (anonymous):

I believe C

OpenStudy (***[isuru]***):

hey...... @KeithAfasCalcLover has made a mistake in his last step.... It should be corrected and the u can have C as answer

OpenStudy (***[isuru]***):

but his steps r correct...!!!

OpenStudy (anonymous):

Oh yeahh haha k thank you isuru I didn't notice that haha

OpenStudy (anonymous):

so its c? i appreciate the help. can you guys help me with another one

OpenStudy (***[isuru]***):

u r welcome

OpenStudy (anonymous):

\[ \eqalign{ \sqrt{18}+\sqrt{48}-\sqrt{50}&=\sqrt{9\times2}-\sqrt{25\times2}+\sqrt{48} \\ &=3\sqrt{2}-5\sqrt{2}+\sqrt{48} \\ &=\sqrt{48}-2\sqrt{2} \\ &=2\sqrt{12}-2\sqrt{2} \\ &=2(\sqrt{12}-\sqrt{2}) \\ &=2(\sqrt{6}\sqrt{2}-\sqrt{2}) \\ &=2[\sqrt{2}(\sqrt{6}-1)] \\ &=(2)^{1.5}(\sqrt{6}-1) \\ &=(2)^{1.5}(\sqrt{2}\sqrt{3}-1) \\ &=2^2\sqrt{3}-2^{1.5} \\ &=4\sqrt{3}-2\sqrt{2} } \] Is that better?

OpenStudy (anonymous):

yeah thanks man. do yo mind helping me with this one?

OpenStudy (***[isuru]***):

fine! by me lol :-)

OpenStudy (anonymous):

294xsquare root of 10 42xsquare root of 10 12x2square root of 10 42square root of 10 times x squared

OpenStudy (anonymous):

these are the answers to the new problem